# Bootstrapping Estimators based on the Block Maxima Method

### v1

@@ 9 Proofs
$\mathbb{P}(\sigma_{r_{n}}\in I_{n,\beta})$ $=\mathbb{P}\Big{[}\sqrt{\frac{n}{r}}\cdot\sigma_{r_{n}}^{-1}\{(F_{\hat{\sigma}_{n}^{*}})^{-1}(\tfrac{\beta}{2})-\hat{\sigma}_{n}\}\leq\sqrt{\frac{n}{r}}\Big{(}\frac{\hat{\sigma}_{n}}{\sigma_{r_{n}}}-1\Big{)}\leq\sqrt{\frac{n}{r}}\cdot\sigma_{r_{n}}^{-1}\{(F_{\hat{\sigma}_{n}^{*}})^{-1}(1-\tfrac{\beta}{2})-\hat{\sigma}_{n}\}\Big{]}$
$=\mathbb{P}\Big{[}(F_{S_{n}^{*}})^{-1}(\tfrac{\beta}{2})\leq S_{n}\leq(F_{S_{n}^{*}})^{-1}(1-\tfrac{\beta}{2})\Big{]},$
the assertion follows from Lemma 4.2 in Bücher and Kojadinovic, (2019). ∎
10 Additional conditions
In order to integrate $\bm{h}(\bm{Z}_{r,i})$ to the limit, we need mild asymptotic integrability conditions. For ${\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}}\}$ the condition is standard. As the circmax-sample permutates the underlying time series on a non-negligible part of the sample, we need a more involved assumption on asymptotic integrability. In most cases both conditions can be verified by similar arguments. Note that (b) implies (a) by letting $t=r$ in the supremum. Moreover, for $k=1$ (disjoint blocks case), (a) and (b) are equivalent.
Condition 10.1 (Asymptotic Integrability).
  1. (a)
  2. (b)
Often we additionally impose that $\nu>2/\omega$ with $\omega$ from Condition 2.2.
Condition 10.2 (Disjoint and sliding blocks bias).
Condition 10.3.
Note in passing that the asymptotic bias $\bm{E}_{\bm{h}}$ already appeared in Bücher and Zanger, (2023) in a slightly modified form in the setting of observing piecewise-stationary time series.
The following result implies that in many situations the asymptotic bias terms $D_{\bm{h},k}$ and $E_{\bm{h}}$ are negligible. The result should be understood as a proof of concept in the sense that one might formulate sharper results involving stronger integrability assumptions (e.g., boundedness of $h_{j}$) or a faster decay of mixing coefficients. Recall the beta-mixing coefficient for stationary time series, $\beta(\ell):=\beta(\sigma(\ldots,\bm{X}_{1},\bm{X}_{0}),\sigma(\bm{X}_{1+\ell},\bm{X}_{2+\ell},\ldots))$ with $\ell\in\mathbb{N}$, where $\beta(\mathcal{B},\mathcal{F})$ denotes the beta dependence coefficient between two sigma-fields $\mathcal{B}$ and $\mathcal{F}$ (Bradley, , 2005).
Lemma 10.4.
Fix $k\in\mathbb{N}_{\geq 2}$ and suppose that Condition 2.1 and 2.2 are met. Then $\bm{D}_{\bm{h},k}=\bm{E}_{\bm{h}}=0$, provided the time series $(\bm{X}_{t})_{t}$ is exponentially $\beta$-mixing (i.e., there exists $c>0,\lambda\in(0,1)$ such that $\beta(\ell)\leq c\lambda^{\ell}$ for all $\ell\in\mathbb{N}$), that $\log n=o(r^{1/2})$ and that Condition 10.1(b) and
$\limsup_{n\to\infty}\sup_{t=1,\ldots,r}\operatorname{E}\Big{[}\Big{\|}h\Big{(}\frac{\max(\bm{X}_{1},\dots,\bm{X}_{t},\tilde{\bm{X}}_{1},\dots,\tilde{\bm{X}}_{r-t})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}\Big{\|}^{2+\nu}\Big{]}<\infty$ (10.3)
$\limsup_{n\to\infty}\sup_{t=\lfloor r^{1/2}\rfloor,\ldots,r}\operatorname{E}\Big{[}\Big{\|}h\Big{(}\frac{\max(\bm{X}_{1},\dots,\bm{X}_{t-\lfloor r^{1/2}\rfloor},\tilde{\bm{X}}_{1},\dots,\tilde{\bm{X}}_{r-t})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}\Big{\|}^{2+\nu}\Big{]}<\infty$ (10.4)
$\limsup_{n\to\infty}\sup_{t=\lfloor r^{1/2}\rfloor,\ldots,r}\operatorname{E}\Big{[}\Big{\|}h\Big{(}\frac{\max(\bm{X}_{1},\dots,\bm{X}_{t-\lfloor r^{1/2}\rfloor},\bm{X}_{t+1},\ldots,\bm{X}_{r})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}\Big{\|}^{2+\nu}\Big{]}<\infty$ (10.5)
Proof.
Note that showing $\sup_{t=1,\ldots,r-1}(n/r)^{1/2}|\operatorname{E}[D_{n,h,k,t}]|=o(1)$ implies $D_{h,k}=0.$ By applying Berbee’s coupling Lemma (Berbee, , 1979) for each fixed $t\in\{1,\ldots,r-1\}$ to the vectors $(\bm{X}_{1},\ldots,\bm{X}_{t})$ and $(\bm{X}_{(k-1)r+t+1},\ldots,\bm{X}_{kr})$ we may assume that the random vector $(\tilde{\bm{X}}_{1},\ldots,\tilde{\bm{X}}_{r-t})$ satisfies $\mathbb{P}\big{(}(\tilde{\bm{X}}_{1},\ldots,\tilde{\bm{X}}_{r-t})\neq(\bm{X}_{(k-1)r+t+1},\ldots,\bm{X}_{kr})\big{)}\leq\beta(r)\leq c\lambda^{r};$ where the last inequality follows by assumption. Thus, by Hölder’s inequality and (10.3),
It remains to show that $E_{h}=0$. Writing
$\hskip 142.26378pt-h\Big{(}\frac{\max(\bm{X}_{1},\ldots,\bm{X}_{t-\ell},\tilde{\bm{X}}_{1},\ldots,\tilde{\bm{X}}_{r-t})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}\Big{]}\Big{|}=o(1)$ (10.6)
$\hskip 142.26378pt-h\Big{(}\frac{\max(\bm{X}_{1},\ldots,\bm{X}_{t-\ell},{\bm{X}}_{t+1},\ldots,{\bm{X}}_{r})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}\Big{]}\Big{|}=o(1)$ (10.7)
$\sup_{t=r/2,\dots,r}\Big{|}\sqrt{\frac{n}{r}}\operatorname{E}\Big{[}h\Big{(}\frac{\max(\bm{X}_{1},\ldots,\bm{X}_{t-\ell},\bm{X}_{t+1},\ldots,\bm{X}_{r})-\bm{b}_{r}}{\bm{a}_{r}}\Big{)}-h(\bm{Z}_{r,1})\Big{]}\Big{|}=o(1).$ (10.8)
The proof of (10.7) is similar to the proof of $D_{h,k}=0$ given above (invoking (10.4) instead of (10.3)), and is therefore omitted (the final argument requires $\log n=o(r^{1/2})$, which is then exactly met by assumption). The proof of (10.7) is a simplified version of the proof of (10.8), so we only prove the latter for the sake of brevity. In view of Condition 10.1(b) and (10.5), an application of Hölder’s inequality implies
[…]
By the union-bound, it is sufficient to show that
$=o(1).$ (10.9)
[…]
We may now successively apply Berbee’s coupling lemma (Berbee, , 1979) to construct a vector $(\check{X}_{j})_{j\in J}$ with iid coordinates that is independent of $X_{i}$ and satisfies $\check{X}_{j}=_{d}X_{j}$ for all $j\in J$ such that
In view of our choice of $\ell=\lfloor r^{1/2}\rfloor$, and letting $\zeta=\min(2,1/\lambda)>1$, the right hand side is bounded by $2r\zeta^{-\sqrt{r}}$, whence $R_{n}$ from (10) can be bounded by
by assumption on $r$. This proves (10), and the proof is finished. ∎
11 Auxiliary results
Proposition 11.1 (Joint weak convergence of circular block maxima).
@@ 11 Auxiliary results
$\lim_{n\to\infty}\mathbb{P}\Big{(}\bm{Z}_{r,(i-1)kr+1+\lfloor\xi r\rfloor}^{({\operatorname{cb}})}\leq\bm{x},\bm{Z}_{r,(i^{\prime}-1)kr+1+\lfloor\xi^{\prime}r\rfloor}^{({\operatorname{cb}})}\leq\bm{y}\Big{)}=G_{1}(\bm{x},\bm{y})=G(\bm{x})G(\bm{y}).$ (11.2)
@@
We start by showing marginal convergence (i.e., $\bm{y}=\bm{\infty}$); the result then corresponds to Proposition 3.1. The case $\xi\leq k-1$ is trivial. For $\xi>k-1$ and fixed $\bm{x}\in\mathbb{R}^{d}$, we may proceed as in the proof of Lemma 2.4 in Bücher and Zanger, (2023) to obtain
$\mathbb{P}\Big{(}\bm{Z}_{r,1+\lfloor\xi r\rfloor}^{({\operatorname{cb}})}\leq\bm{x}\Big{)}$ $=\mathbb{P}\Big{(}\frac{\max(\bm{X}_{1+\lfloor\xi r\rfloor},\ldots,\bm{X}_{kr})-\bm{b}_{r}}{\bm{a}_{r}}\leq\bm{x},\frac{\max(\bm{X}_{1},\ldots,\bm{X}_{r+\lfloor\xi r\rfloor-kr})-\bm{b}_{r}}{\bm{a}_{r}}\leq\bm{x}\Big{)}$
@@
Arguing as in the proof of Lemma 8.9 from Bücher and Staud, (2024), Condition 2.1 and the previous three displays yield, for $n\to\infty,$
$\mathbb{P}(\bm{Z}_{r,1+\lfloor\xi r\rfloor}^{({\operatorname{cb}})}\leq\bm{x})$ $=G^{k-\xi}(\bm{x})G^{1+\xi-k}(\bm{x})+o(1)=G(\bm{x})+o(1).$
We proceed by proving (11.1), and start by considering the case $|\xi-\xi^{\prime}|>k-1$. Without loss of generality let $\xi<\xi^{\prime}$ such that $1+\lfloor\xi r\rfloor\leq r$ and $1+\lfloor\xi^{\prime}r\rfloor>(k-1)r$. Then, by the same arguments as for the marginal convergence,
@@
Next, consider the case $|\xi-\xi^{\prime}|\leq k-1$, and again assume $\xi<\xi^{\prime}$. There are two subcases to handle; first let $\xi^{\prime}\in(k-1,k].$ In that case we have $\xi\geq 1,$ hence there is no overlapping from the left (induced by the $\operatorname{circ{\text{-}}max}$-operation). We may then proceed in a similar (but simpler) way as for $|\xi-\xi^{\prime}|>k-1$ to obtain the stated limit. Alternatively, we have $\xi^{\prime}\leq k-1.$ In that case, we are in the sliding block maxima case, where the result is known, see e.g. Lemma 8.6 in Bücher and Staud, (2024).
Finally, the assertion in (11.2) follows from similar arguments as in Lemmas A.7, A.8 from Bücher and Segers, 2018b , which yield
$\mathbb{P}\big{(}\bm{Z}_{(i-1)kr+1+\lfloor\xi r\rfloor}^{({\operatorname{cb}})}\leq\bm{x},\bm{Z}_{(i^{\prime}-1)kr+1+\lfloor\xi^{\prime}r\rfloor}^{({\operatorname{cb}})}\leq\bm{y}\big{)}$ $=\mathbb{P}\big{(}\bm{Z}_{1+\lfloor\xi r\rfloor}^{({\operatorname{cb}})}\leq\bm{x}\big{)}\cdot\mathbb{P}\big{(}\bm{Z}_{1+\lfloor\xi^{\prime}r\rfloor}^{({\operatorname{cb}})}\leq\bm{y}\big{)}+o(1)$
@@
Lemma 11.2 (Multiplier central limit theorem).
Suppose Conditions 2.1 and 2.2 are met. Fix $k\in\mathbb{N}$ and let $\bm{h}=(h_{1},\dots,h_{q})^{\top}$ satisfy Condition 10.1 with $\nu>2/\omega$ and $\omega$ from Condition 2.2. Let $Y_{n,1},\ldots,Y_{n,m(k)}$ be a triangular array of rowwise independent and identically distributed random variables that is independent of $(\bm{X}_{1},\bm{X}_{2},\dots)$ such that $\limsup_{n\to\infty}\operatorname{E}[|Y_{n,1}|^{2+\nu}]<\infty$ with $\nu$ from Condition 10.1 and such that the limit $\mu_{2}:=\lim_{n\to\infty}\operatorname{E}[Y_{n,1}^{2}]>0$ exists. Then, for $n\to\infty$ and with $\Sigma_{\bm{h}}^{({\operatorname{mb}})}$ from (2.4), we have
$\bar{\mathbb{G}}_{n,r}^{\circ}\bm{h}:=\sqrt{\frac{n}{r}}\frac{1}{n}\sum_{i=1}^{m(k)}Y_{n,i}\sum_{s\in I_{kr,i}}\big{\{}\bm{h}(\bm{Z}_{r,s}^{({\operatorname{cb}})})-\operatorname{E}[\bm{h}(\bm{Z}_{r,s}^{({\operatorname{cb}})})]\big{\}}\rightsquigarrow\begin{cases}\mathcal{N}_{q}(0,\mu_{2}\Sigma_{\bm{h}}^{({\operatorname{db}})})&\text{ if }k=1,\\ \mathcal{N}_{q}(0,\mu_{2}\Sigma_{\bm{h}}^{({\operatorname{sb}})})&\text{ if }k\in\{2,3,\dots\}.\end{cases}$
@@
We only consider the case $k\geq 2$; the case $k=1$ is a simple modification. By the Cramér-Wold Theorem, it is sufficient to consider the case $q=1$. Since $m(k)=n/(kr)$, we may write $\bar{\mathbb{G}}_{n,r}^{\circ}h=\sqrt{\frac{n}{r}}\frac{1}{m(k)}\sum_{i=1}^{m(k)}Y_{n,i}D_{n,i}$, with $D_{n,i}=\frac{1}{kr}\sum_{s\in I_{kr,i}}\{h(\bm{Z}_{r,s}^{({\operatorname{cb}})})-\operatorname{E}[h(\bm{Z}_{r,s}^{({\operatorname{cb}})})]\}$ from (9.2). Then, since $(Y_{n,i}D_{n,i})_{i}$ is uncorrelated,
$\operatorname{Var}(\bar{\mathbb{G}}_{n,r}^{\circ}h)$ $=k\operatorname{E}[Y_{n,1}^{2}]\operatorname{Var}(D_{n,1})=\mu_{2}\Sigma_{h}^{({\operatorname{sb}})}+o(1)$
by (9.4). This implies the assertion for $\Sigma_{h}^{({\operatorname{sb}})}=0$.
If $\Sigma_{h}^{({\operatorname{sb}})}>0$, we may follow a similar line of reasoning as in the proof of Theorem 3.2: decompose
@@
The next result is possibly well-known; we skip its elementary proof (see also Section 2 in Klaassen and Wellner, , 1992).
Lemma 11.3 (Poissonization).
@@
$\hskip 28.45274pt+\operatorname{E}[h(U)h(W)G^{1-\xi}(U\wedge W)]+\operatorname{E}[h^{2}(V)G^{2\xi}(V)]\equiv 2I_{1}+I_{2}+I_{3}.$ (11.3)
@@
Similarly, we have $I_{2}=\xi^{2}(1-\xi)\int_{-\infty}^{\infty}G^{\xi-2}(x)g(x)H^{2}(x)\,\mathrm{d}x\geq 0.$ In view of (11) and $I_{3}\geq 0$ this proves the first assertion.
For the second assertion note that $I_{3}=(1-\xi)\operatorname{E}[h^{2}(\bm{Z})G^{\xi}(\bm{Z})]\leq\operatorname{Var}(h(\bm{Z}))<\infty$ and $I_{j}\geq 0$ for $j=1,2,3$. By dominated convergence and (11) we then have
@@
11.1 A general result on bootstrapping the Fréchet independence MLE

### v2

@@ Appendix B Proofs
$\mathbb{P}(\sigma_{r_{n}}\in I_{n,\sigma}^{[{\operatorname{s}},{\operatorname{c}}]})$ $=\mathbb{P}\Big{[}\sqrt{\frac{n}{r}}\cdot\sigma_{r_{n}}^{-1}\{(F_{\hat{\sigma}_{n}^{[{\operatorname{c}}],*}})^{-1}(\tfrac{\beta}{2})-\hat{\sigma}_{n}^{[{\operatorname{c}}]}\}\leq\sqrt{\frac{n}{r}}\Big{(}\frac{\hat{\sigma}_{n}^{[{\operatorname{s}}]}}{\sigma_{r_{n}}}-1\Big{)}\leq\sqrt{\frac{n}{r}}\cdot\sigma_{r_{n}}^{-1}\{(F_{\hat{\sigma}_{n}^{[{\operatorname{c}}],*}})^{-1}(1-\tfrac{\beta}{2})-\hat{\sigma}_{n}^{[{\operatorname{c}}]}\}\Big{]}$
$=\mathbb{P}\Big{[}(F_{S_{n}^{[{\operatorname{c}}],*}})^{-1}(\tfrac{\beta}{2})\leq S_{n}^{[{\operatorname{s}}]}\leq(F_{S_{n}^{[{\operatorname{c}}],*}})^{-1}(1-\tfrac{\beta}{2})\Big{]},$
the assertion follows from Lemma 4.2 in BucKoj2019note. ∎
Appendix C Auxiliary results
Proposition C.1 (Joint weak convergence of circular block maxima).
@@ Appendix C Auxiliary results
$\lim_{n\to\infty}\mathbb{P}\Big{(}\bm{Z}_{r,(i-1)kr+1+\lfloor\xi r\rfloor}^{[{\operatorname{c}}]}\leq\bm{x},\bm{Z}_{r,(i^{\prime}-1)kr+1+\lfloor\xi^{\prime}r\rfloor}^{[{\operatorname{c}}]}\leq\bm{y}\Big{)}=G_{1}(\bm{x},\bm{y})=G(\bm{x})G(\bm{y}).$ (C.2)
@@
We start by showing marginal convergence (i.e., $\bm{y}=\bm{\infty}$); the result then corresponds to Proposition 3.1. The case $\xi\leq k-1$ is trivial. For $\xi>k-1$ and fixed $\bm{x}\in\mathbb{R}^{d}$, we may proceed as in the proof of Lemma 2.4 in BucZan23 to obtain
$\mathbb{P}\Big{(}\bm{Z}_{r,1+\lfloor\xi r\rfloor}^{[{\operatorname{c}}]}\leq\bm{x}\Big{)}$ $=\mathbb{P}\Big{(}\frac{\max(\bm{X}_{1+\lfloor\xi r\rfloor},\ldots,\bm{X}_{kr})-\bm{b}_{r}}{\bm{a}_{r}}\leq\bm{x},\frac{\max(\bm{X}_{1},\ldots,\bm{X}_{r+\lfloor\xi r\rfloor-kr})-\bm{b}_{r}}{\bm{a}_{r}}\leq\bm{x}\Big{)}$
@@
Arguing as in the proof of Lemma 8.9 from BucSta24, Condition 2.1 and the previous three displays yield, for $n\to\infty,$
$\mathbb{P}(\bm{Z}_{r,1+\lfloor\xi r\rfloor}^{[{\operatorname{c}}]}\leq\bm{x})$ $=G^{k-\xi}(\bm{x})G^{1+\xi-k}(\bm{x})+o(1)=G(\bm{x})+o(1).$
We proceed by proving (C.1), and start by considering the case $|\xi-\xi^{\prime}|>k-1$. Without loss of generality let $\xi<\xi^{\prime}$ such that $1+\lfloor\xi r\rfloor\leq r$ and $1+\lfloor\xi^{\prime}r\rfloor>(k-1)r$. Then, by the same arguments as for the marginal convergence,
@@
Next, consider the case $|\xi-\xi^{\prime}|\leq k-1$, and again assume $\xi<\xi^{\prime}$. There are two subcases to handle; first let $\xi^{\prime}\in(k-1,k].$ In that case we have $\xi\geq 1,$ hence there is no overlapping from the left (induced by the $\operatorname{circ{\text{-}}max}$-operation). We may then proceed in a similar (but simpler) way as for $|\xi-\xi^{\prime}|>k-1$ to obtain the stated limit. Alternatively, we have $\xi^{\prime}\leq k-1.$ In that case, we are in the sliding block maxima case, where the result is known, see e.g. Lemma 8.6 in BucSta24.
Finally, the assertion in (C.2) follows from similar arguments as in Lemmas A.7, A.8 from BucSeg18a, which yield
$\mathbb{P}\big{(}\bm{Z}_{(i-1)kr+1+\lfloor\xi r\rfloor}^{[{\operatorname{c}}]}\leq\bm{x},\bm{Z}_{(i^{\prime}-1)kr+1+\lfloor\xi^{\prime}r\rfloor}^{[{\operatorname{c}}]}\leq\bm{y}\big{)}$ $=\mathbb{P}\big{(}\bm{Z}_{1+\lfloor\xi r\rfloor}^{[{\operatorname{c}}]}\leq\bm{x}\big{)}\cdot\mathbb{P}\big{(}\bm{Z}_{1+\lfloor\xi^{\prime}r\rfloor}^{[{\operatorname{c}}]}\leq\bm{y}\big{)}+o(1)$
@@
Lemma C.2 (Multiplier central limit theorem).
Suppose Conditions 2.1 and 2.2 are met. Fix $k\in\mathbb{N}$ and let $\bm{h}=(h_{1},\dots,h_{q})^{\top}$ satisfy Condition 9.1 with $\nu>2/\omega$ and $\omega$ from Condition 2.2. Let $Y_{n,1},\ldots,Y_{n,m(k)}$ be a triangular array of rowwise independent and identically distributed random variables that is independent of $(\bm{X}_{1},\bm{X}_{2},\dots)$ such that $\limsup_{n\to\infty}\operatorname{E}[|Y_{n,1}|^{2+\nu}]<\infty$ with $\nu$ from Condition 9.1 and such that the limit $\mu_{2}:=\lim_{n\to\infty}\operatorname{E}[Y_{n,1}^{2}]>0$ exists. Then, for $n\to\infty$ and with $\Sigma_{\bm{h}}^{[{\operatorname{m}}]}$ from (2.4), we have
$\bar{\mathbb{G}}_{n,r}^{\circ}\bm{h}:=\sqrt{\frac{n}{r}}\frac{1}{n}\sum_{i=1}^{m(k)}Y_{n,i}\sum_{s\in I_{kr,i}}\big{\{}\bm{h}(\bm{Z}_{r,s}^{[{\operatorname{c}}]})-\operatorname{E}[\bm{h}(\bm{Z}_{r,s}^{[{\operatorname{c}}]})]\big{\}}\rightsquigarrow\begin{cases}\mathcal{N}_{q}(0,\mu_{2}\Sigma_{\bm{h}}^{[{\operatorname{d}}]})&\text{ if }k=1,\\ \mathcal{N}_{q}(0,\mu_{2}\Sigma_{\bm{h}}^{[{\operatorname{s}}]})&\text{ if }k\in\{2,3,\dots\}.\end{cases}$
@@
We only consider the case $k\geq 2$; the case $k=1$ is a simple modification. By the Cramér-Wold Theorem, it is sufficient to consider the case $q=1$. Since $m(k)=n/(kr)$, we may write $\bar{\mathbb{G}}_{n,r}^{\circ}h=\sqrt{\frac{n}{r}}\frac{1}{m(k)}\sum_{i=1}^{m(k)}Y_{n,i}D_{n,i}$, with $D_{n,i}=\frac{1}{kr}\sum_{s\in I_{kr,i}}\{h(\bm{Z}_{r,s}^{[{\operatorname{c}}]})-\operatorname{E}[h(\bm{Z}_{r,s}^{[{\operatorname{c}}]})]\}$ from (B.2). Then, since $(Y_{n,i}D_{n,i})_{i}$ is uncorrelated,
$\operatorname{Var}(\bar{\mathbb{G}}_{n,r}^{\circ}h)$ $=k\operatorname{E}[Y_{n,1}^{2}]\operatorname{Var}(D_{n,1})=\mu_{2}\Sigma_{h}^{[{\operatorname{s}}]}+o(1)$
by (B.4). This implies the assertion for $\Sigma_{h}^{[{\operatorname{s}}]}=0$.
If $\Sigma_{h}^{[{\operatorname{s}}]}>0$, we may follow a similar line of reasoning as in the proof of Theorem 3.2: decompose
@@
The next result is possibly well-known; we skip its elementary proof (see also Section 2 in KlaWel91).
Lemma C.3 (Poissonization).
@@
$\hskip 28.45274pt+\operatorname{E}[h(U)h(W)G^{1-\xi}(U\wedge W)]+\operatorname{E}[h^{2}(V)G^{2\xi}(V)]\equiv 2I_{1}+I_{2}+I_{3}.$ (C.3)
@@
Similarly, we have $I_{2}=\xi^{2}(1-\xi)\int_{-\infty}^{\infty}G^{\xi-2}(x)g(x)H^{2}(x)\,\mathrm{d}x\geq 0.$ In view of (C) and $I_{3}\geq 0$ this proves the first assertion.
For the second assertion note that $I_{3}=(1-\xi)\operatorname{E}[h^{2}(\bm{Z})G^{\xi}(\bm{Z})]\leq\operatorname{Var}(h(\bm{Z}))<\infty$ and $I_{j}\geq 0$ for $j=1,2,3$. By dominated convergence and (C) we then have
@@
C.1 A general result on bootstrapping the Fréchet independence MLE